OpenStudy (anonymous):

solve using the addition principle 5=x+4

OpenStudy (ash2326):

We know that the Addition Principle states that adding (or subtracting) the same number to both sides of an equation does not change its solution set.

OpenStudy (ash2326):

Let's use this here we have \[5=x+4\] subtract both sides by 4 \[5-4=x+4-4\] What do you get now @jlamphere?

OpenStudy (anonymous):

1=0+4 right

OpenStudy (ash2326):

Check again this is not correct

OpenStudy (anonymous):

x=1+4-4

OpenStudy (ash2326):

What do you get finally?

OpenStudy (anonymous):

x=5-4 x=1

OpenStudy (ash2326):

Yeah great work:D

OpenStudy (anonymous):

thanks